OpenStudy (cookie_2046):

The length of a rectangle is 5 more than 3 times its width. The perimeter of the rectangle is 58 inches. What is the length of the rectangle? Question 10 options: 23 in. 29 in. 32 in. 35 in.

OpenStudy (campbell_st):

so let the width be w then the length is 3w + 5 to the perimeter is P = 2( l + w) so using the information P = 2(w + 3w + 5) or P = 8w + 10 you know the perimeter so 58 = 8w + 10 solve for w then from there you can find the length of the rectangle